' Part 'A' Trinh is starting a savings plan at the bank. The bank will automatically transfer any amount she wants from her checking account into a savings account. After the first month, the amount transferred each month will be the first month plus '$1250. Write an equation to model this situation. Clearly identify the variables you are using.


Part B Trinh wants to save '$25.' total after 15 months to put towards a vacation. Use the equation you wrote in Part A to determine how much money she needs to transfer the first month to achieve this goal. Show your work or explain how you found your answer.


1
Expert's answer
2021-03-05T07:34:27-0500

A. "x+1250=x"

in the first month


B. x+(x+1250)15-x=25

15x+18750=25

15x=-18725

x=-1248.33

she needs to transfer the first month